UT20090707 LUCIFER Commissioning

Observers: WSeifert, NAgeorges (LBT)
Telescope Operator: D Gonzalez Huerta
Software Support:
Instrument Support:
Telescope Support:

Summary

Details (UT unless otherwise noted)

Notes from the DGH nightly report

01:30 SX raised
02:00 Telescope readied.
02:30 Vent doors open
02:35 Open enclosure
02:46 first preset issued in track mode. Pointing corrections made
02:56 Preset issued in active mode. new target. CCD saturaded (GCS). Sky still
too bright.
03:01 Mirror ventilation started
03:02 Vent doors closed for winds (9.9 m/sec)
03:04 Focus correction applied. -0.2 in Z global
03:05 Preset resent on same field. Guide star found!!
03:11 DIMM started.
03:13 DIMM starts measuring. 0.99
04:22 Primary mirror reaches limits on RX and RY. No actions taken.
04:28 Guiding stopped. Calibration unit started
04:35 preset issued in active mode. slewing to new field.
04:40 Clouds in the east.
05:10 Focus applied on primary mirror, global z 4 = -0.1
05:45 Calibration unit started
05:55 Preset issued in active mode to new field. No guide star found.
06:06 Preset issued in track mode in the same field.
06:11 Focus correction applied, global Z=0.1
06:37 Preset issued in active mode to new field. No guide star found
07:02 Preset issued in active mode, guide star found.
07:54 Preset issued in active mode to new field.It failed, no guide star found
7:58 AZ encoders reset in order to equal the absolute encoders value. Issuetrak 2263
07:59 preset sent in track mode in the same field.
08:04 close for clouds above the telescope. (the unfriendly type of clouds).
08:17 Telescope at the zenith for calibrations.
08:58 Open enclosure. resume observing
09:04 Preset issued in track mode. checking pointing
09:15 DIMM restarted. Seeing = 1.2
09:28 Preset issued in active mode in the same field. No guide star found.
09:34 Pointing corrections
09:38 Preset resent in active mode. Guide star found
09:46 Vent doors open, winds speed ok.
10:24 Guiding stopped. Calibration unit started
10:45 Back to active mode and guiding.
11:33 Vent doors closed for winds (9 m/sec)
12:00 Shut down
12:02 Calibrations started with telescope at the zenith.

-- JohnHill - 08 Jul 2009
I Attachment Action Size Date Who Comment
daq.20090707.gifgif daq.20090707.gif manage 37 K 08 Jul 2009 - 18:04 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r1 - 08 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback